# Evaluating Large Language Models Across 100+ Languages: A Unified Benchmark for Speech and Text Tasks

### v2

@@ Abstract
Large Language models (LLMs) have demonstrated impressive performance on a wide range of tasks, including in multimodal settings such as speech. However, their evaluation is often limited to English and a few high-resource languages. For low-resource languages, there is no standardized evaluation benchmark. In this paper, we address this gap by introducing mSTEB, a new benchmark to evaluate the performance of LLMs on a wide range of tasks covering language identification, text classification, question answering, and translation tasks on both speech and text modalities. We evaluated the performance of leading LLMs such as Gemini 2.0 Flash and GPT-4o (Audio) and state-of-the-art open models such as Qwen 2 Audio and Gemma 3 27B. Our evaluation shows a wide gap in performance between high-resource and low-resource languages, especially for languages spoken in Africa and Americas/Oceania. Our findings show that more investment is needed to address their under-representation in LLMs coverage.
@@ I Introduction
Recently, Large language models have demonstrated impressive performance on various tasks [1, 2, 3], including in multimodal settings such as vision and speech [4, 5, 6, 7, 8]. However, their evaluation is often limited to the English language and a few other high-resource languages, such as Chinese and French. Moreover, many frontier LLMs claim multilingual capabilities but they often fail to disclose the specific languages included in their pre-training data. While extensive evaluations exist for text-based tasks across many languages [9, 10, 11], comparable assessments for speech-based tasks, particularly those covering hundreds of languages, remain a significant gap.
There has been some progress in expanding speech models evaluation to many tasks, however, they often do not cover LLM evaluations and only focus on a few languages [12]. The tasks with broad language coverage are typically language identification (LID) and automatic speech recognition (ASR), particularly when evaluating speech representation models [13, 14]. In contrast, more recent audio LLMs such as SALMONN [6] and Qwen 2 Audio [4] generally address a wide range of tasks but support only a limited set of languages. Some of these tasks are more challenging than the usual LID and ASR tasks that are often evaluated for many low-resource languages; they include audio captioning, scene classification, audio question answering, and speech-to-text translation. Therefore, for a more realistic evaluation of LLMs as they become more multilingual, it is important to evaluate them on a wide range of tasks and a diverse set of languages including low-resource languages.
@@
Leveraging mSTEB, we evaluated five state-of-the art LLMs: GPT-4oGPT-4o Audio,  Gemini 2.0 Flash (for both modalities), Qwen 2 Audio, and Gemma 3 27B. The last two models are open-weight models while the others are proprietary. Our evaluation shows a large performance gap between the open models and proprietary models for both texts and speech tasks, however, the gap is much wider for the speech-based tasks, where Qwen 2 Audio is worse on all languages. In contrast, in the text-based evaluations, we found Gemma 3 27B to be very competitive to the proprietary models on European and East Asian languages (CJK). Furthermore, there is a wide gap for the regions with a lot of low-resource languages such as Africa and Americas/Oceania regions. Our analysis demonstrates large gap in performance for many low-resource languages compared to high-resource languages, which shows that more efforts are needed to address their under-representation in LLMs.
@@ IV Experimental setup
LID You are an assistant specialized in identifying spoken languages. Please identify the language in the audio clip below. Return only the name of the language, no other text.11footnotemark: 1
@@
Baseline models: For each of the task, we evaluate the performance of an existing supervised model, if available. When there is no training data, we make use of cross-lingual transfer results as the baseline by fine-tuning on English training data, and evaluating on the remaining languages. For the speech-based tasks, we make use of MMS-LID-2048 (trained on 2048 languages) [41] for LID, SeamlessM4T v2 [42] for the ASR and S2TT tasks, and for the TC and RC-QA, we make use of the baseline from Fleurs-SLU paper [24] which is based on a cascaded approach—where audio utterances are first transcribed by an ASR/S2TT model (Seamless, followed by an evaluation on a text classifier (similar to Translate-Test). [^2] This involves machine (or speech) translating test set from another language to English, followed by evaluating an English-only model. Similarly, for the text-based tasks, we evaluate GlotLID [43] trained on over 2,000 languages on LID, NLLB-3.3B [21] on machine translation, Translate-Train [^3] English training data machine translated to other languages, then combined together for fine-tuning the base model based on XLM-V [44] for RC-QA, and cross-lingual transfer of English model to other non-English languages for TC and NLI. For the TC, we make use of the NLLB-LLM2Vec encoder [45]—a model that stacks the encoder part of NLLB with the LLaMa-3 8B [46] decoder model to create an impressive multilingual sentence transformer model, based on LLM2Vec [47], while for NLI, we leveraged mDeBERTaV3 [48]. For both TC, RC-QA, we make use of the English training data provided by the authors of the dataset, while for NLI, we fine-tuned the English MNLI [49] dataset.
@@
Evaluation tools and hardware requirements: We used Open API for the GPT-4o and Google Cloud API for Gemini-2.0 Flash models, to prompt and receive responses. For Gemma 3 27B, we used two NVIDIA L40S 48GB GPUs and served it via vLLM [53] for faster inference. However, for the Qwen 2 Audio, we used two NVIDIA GeForce RTX 3090 GPUs for LID and topic classification, and Quadro RTX 6000 and a Tesla V100S-PCIE-32GB for speech translation, RC-QA, and ASR.
@@ V Results And Discussions
Africa 0.0 56.6 68.5 28.2 55.3 68.5 26.3 55.5 83.9 60.1 18.0 13.2 37.6 18.2 64.2
Americas/Oceania 0.3 98.1 96.7 15.2 66.9 44.7 27.1 50.9 146.2 31.7 19.5 11.9 27.2 14.2 64.1
Asia (S) 4.5 91.8 97.5 32.8 79.8 83.0 30.3 71.5 113.7 35.1 7.7 16.0 58.3 22.5 84.0
Asia (SE) 10.5 88.8 91.0 33.0 76.0 82.3 31.4 75.8 97.8 35.0 7.3 16.0 57.8 24.9 83.0
Asia (W, C) 21.7 80.2 88.0 28.4 79.1 82.1 30.1 74.7 111.7 27.2 6.7 15.6 55.5 26.7 81.6
Asia (E) 59.2 90.5 98.1 58.7 80.5 83.5 43.5 77.3 48.2 23.2 13.9 34.0 52.8 53.8 86.3
Europe (W, N, S) 28.6 86.1 84.5 54.5 84.0 82.1 41.6 81.7 62.6 12.4 7.3 29.6 58.3 41.5 82.8
Europe (E) 11.8 90.4 94.1 53.6 85.6 84.9 36.6 82.1 98.2 12.0 9.9 23.9 62.2 34.0 87.0
Average (macro) 17.1 85.3 89.8 38.0 75.9 76.4 33.4 71.2 95.3 29.6 11.3 20.0 51.2 29.5 79.1
Average (micro) 15.8 81.2 86.1 40.4 76.0 79.5 33.3 72.6 88.4 30.1 10.3 20.5 53.5
@@ V-A Overall Results
Figure 1: Average results by tasks. For the ASR, we reported “$100-CER$” value.
@@ V-A1 Results by Task
Figure 1 shows the result of various LLMs and the best result obtained from a task-specific model (i.e. Task Baseline). For the speech tasks, we find that for classification tasks (i.e. LID, TC, and RC-QA), our result shows that the task baseline achieved much better result, for example, the MMS-LID accuracy is more than $+10$ points better than the best LLM (Gemini 2.0 Flash). Similarly, for TC and RC-QA, the performance gap is more than $+8$ and $+13$ points respectively. Overall, we saw a wide performance gap between Qwen 2 Audio and Gemini 2.0 Flash which clearly shows that the former only covers a few languages. For more challenging tasks such as ASR and S2TT, surprisingly, we found Gemini 2.0 Flash to be better than SeamlessM4T v2, one of the SOTA task-specific baseline for both tasks. This result shows that Gemini 2.0 Flash should be preferred for these speech tasks especially for low-resource languages rather than their task-specific variants. For the text-based tasks, the finding is different, we find that the gap in performance between an open weight model such as Gemma 3 27B and Gemini 2.0 Flash or GPT-4o is much smaller. Similarly, the performance of task baselines are also very similar to those of the LLMs on almost all tasks except for LID and RC-QA. For LID, task baselines (GlotLID) is better by $+5$ points accuracy when compared to Gemini 2.0 Flash, while for RC-QA, all LLMs are significantly better than task baseline based on Translate-TRAIN (on XLM-V) with over $+17$ points.
@@ V-A2 Results by Region
Table  III and IV shows the result across eight different geographical groups. For the speech tasks, we find that the African and the Americas/Oceania region achieved the worst result across all the covered regions probably because most of the low-resource languages are concentrated there. While Qwen 2 Audio model achieved the worst results among the LLMs we evaluated, we find that its performance is slightly better for Asia (E) and Europe (W, N, S) which may reflect where most of the models pre-training data are coming from. For example on ASR and S2TT, it achieved less than 50 CER and more than 30 ChrF on East Asia region. Gemini 2.0 Flash on the other hand had less than 15 CER for all regions except Americas/Oceania and Africa. However, GPT-4o was worse than Gemini 2.0 Flash across all regions we tested, it only achieved less than 15 CER for the European regions. The worst performance of GPT-4o is also for African languages at 60.1 CER. We observe similar trend for S2TT where Gemini 2.0 Flash achieved more than 50 ChrF++ on all regions except the Americas/Oceania and African regions. For classification tasks, Gemini 2.0 Flash had over 70 accuracy points on TC and RC-QA except for the African and Americas region; however, for the LID, the most language confusion is for the African region. For the text-based tasks, the findings are similar across regions, but most regions have high performance except the African region. While we observe the Americas region also has high performance across tasks including MT, this is not true for NLI where it achieved the lowest ($46.8$), more than $-17$ points behind Africa. One potential reason is that large evaluation datasets such as Flores and SIB-200 only cover few Americas/Oceania languages (about 5, including two Creole languages), however AmericasNLI has more truly low-resource languages (e.g. Otomi and Wixarika in Mexico), which provides more realistic resource-level and performance of Americas languages.
@@ V-A3 Results by Language Family
Text-based tasks achieve overall higher performance: In general, performance on texts is higher than performance on speech for same tasks. This implies that current LLMs are better on texts than speech, which implies that more investment in developing better audio LLMs are needed. On the LID task, several language families especially those that use a different script have perfect accuracy score using Gemini 2.0 Flash, however, we noticed some mis-classifications for the spoken utterances. Similar trend is observed for TC and RC-QA classification tasks; on TC, all language families except Nilotic achieved more than $80$ accuracy point. Also, for the translation tasks, several language families achieved impressive ChrF++ of $60$ points and above in xx-en direction while only one language family (Uralic) achieved up to 60 ChrF++ in speech translation.
Language families in Africa achieved the lowest performance: For both text and speech modalities, the language families with the lowest performance are: Atlantic-Congo, Nilotic, and Afro-Asiatic, many or all of the languages in the first two language families are often categorized as low-resourced according to Joshi’s classification [57] and are only located in Africa. This confirms our findings in Sec. V-A2. Apart from having low performance, the gap in performance is also quite wide. For example, the average ChrF++ for S2TT in Nilotic and Atlantic-Congo is $-36.5$ and $-25.9$ lower when compared to Indo-European languages. Similarly, for RC-QA, the performance of Indo-European is $78.4$ while Atlantic-Congo, Nilotic, and Afro-Asiatic achieved $-36.5$, $-25.9$, and $-25.9$ respectively.
@@ V-B Detailed analysis of task specific results
The Global NLI dataset has the highest representation of Americas/Oceania amongst all the other datasets that we evaluated, with 11 out of the 59 languages being from this region. This is due to the Americas NLI dataset [30] and only 3 languages of this region are present in the other text-based tasks. On this task, GPT-4o performs the best overall and in most regions. Americas/Oceania is the region with the lowest performance in this task. We see a performance gap of $-35$ points between Americas/Oceania (49.1%) and Europe (W, N, S) (84.8%), which is larger than that present in other tasks. The NLI results show that many Americas/Oceanian languages are less covered by many LLMs, although only few multilingual datasets includes these indigeneous languages. More efforts are needed to expand benchmarks for this region of languages.
@@ Acknowledgment
This research was supported by the Google grant via Mila. We would also like to thank Google Cloud for the GCP credits Award through the Gemma 2 Academic Program, and OpenAI for providing us access for providing API credits through their Researcher Access API Program.

### v3

@@ Abstract
Large Language models (LLMs) have demonstrated impressive performance on a wide range of tasks, including in multimodal settings such as speech. However, their evaluation is often limited to English and a few high-resource languages. For low-resource languages, there is no standardized evaluation benchmark. In this paper, we address this gap by introducing mSTEB, a new benchmark to evaluate the performance of LLMs on a wide range of tasks covering language identification, text classification, question answering, and translation tasks on both speech and text modalities. We evaluated the performance of leading LLMs such as Gemini 2.0 Flash and GPT-4o and state-of-the-art open models such as Qwen 2 Audio and Gemma 3 27B. Our evaluation shows a wide gap in performance between high-resource and low-resource languages, especially for languages spoken in Africa and Americas/Oceania. Our findings show that more investment is needed to address their under-representation in LLMs coverage.
@@ I Introduction
Recently, Large language models (LLMs) have demonstrated impressive performance on various tasks [1, 2, 3], including in multimodal settings such as vision and speech [4, 5, 6, 7, 8]. However, their evaluation is often limited to the English language and a few other high-resource languages, such as Chinese and French. Many frontier LLMs claim multilingual capabilities yet rarely disclose the languages in their pre-training data. While extensive evaluations exist for text-based tasks across many languages  [9, 10, 11], comparable assessments for speech-based tasks, especially across hundreds of languages, remain scarce.
While speech model evaluations have expanded to many tasks, they often exclude LLMs and focus on only a few languages [12]. Broad language coverage is mostly seen in language identification (LID) and automatic speech recognition (ASR), especially for speech representation models [13, 14]. In contrast, recent audio LLMs like SALMONN [6] and Qwen 2 Audio [4] handle diverse tasks but support few languages. Some of these tasks such as audio captioning, scene classification, audio QA, and speech-to-text translation—are often more challenging than LID and ASR, especially for low-resource languages. Realistic multilingual LLM evaluation thus requires a broad range of tasks and languages, including low-resource ones.
@@
Leveraging mSTEB, we evaluated five state-of-the art LLMs: GPT-4o, GPT-4o Audio,  Gemini 2.0 Flash (for both modalities), Qwen 2 Audio, and Gemma 3 27B. The last two models are open-weight models while the others are proprietary. Our evaluation shows a large performance gap between the open models and proprietary models for both texts and speech tasks, however, the gap is much wider for the speech-based tasks, where Qwen 2 Audio is worse on all languages. In contrast, in the text-based evaluations, we found Gemma 3 27B to be very competitive to the proprietary models on European and East Asian languages (CJK). Furthermore, there is a wide gap for the regions with a lot of low-resource languages such as Africa and Americas/Oceania regions. Our analysis demonstrates large gap in performance for many low-resource languages compared to high-resource languages, which shows that more efforts are needed to address their under-representation in LLMs. We release our code, data and leaderboard on GitHub and HuggingFace platform.  [^1] The code and datasets are available on https://github.com/McGill-NLP/mSTEB. We also release a leaderboard : https://huggingface.co/spaces/McGill-NLP/msteb_leaderboard
@@ IV Experimental setup
LID You are an assistant specialized in identifying spoken languages. Please identify the language in the audio clip below. Return only the name of the language, no other text.22footnotemark: 2
@@
Baseline models: For each of the task, we evaluate the performance of an existing supervised model, if available. When there is no training data, we make use of cross-lingual transfer results as the baseline by fine-tuning on English training data, and evaluating on the remaining languages. For the speech-based tasks, we make use of MMS-LID-2048 (trained on 2048 languages) [41] for LID, SeamlessM4T v2 [42] for the ASR and S2TT tasks, and for the TC and RC-QA, we make use of the baseline from Fleurs-SLU paper [24] which is based on a cascaded approach—where audio utterances are first transcribed by an ASR/S2TT model (Seamless), followed by an evaluation on a text classifier (similar to Translate-Test).Similarly, for the text-based tasks, we evaluate GlotLID [43] trained on over 2,000 languages on LID, NLLB-3.3B [21] on machine translation, Translate-Trainbased on XLM-V [44] for RC-QA, and cross-lingual transfer of English model to other non-English languages for TC and NLI. For the TC, we make use of the NLLB-LLM2Vec encoder [45]—a model that stacks the encoder part of NLLB with the LLaMa-3 8B [46] decoder model to create an impressive multilingual sentence transformer model, based on LLM2Vec [47], while for NLI, we leveraged mDeBERTaV3 [48]. For both TC, RC-QA, we make use of the English training data provided by the authors of the dataset, while for NLI, we fine-tuned the English MNLI [49] dataset.
@@
Evaluation tools and hardware requirements: We used the OpenAI API for GPT-4o (Audio) and Google Cloud API for Gemini-2.0 Flash. Gemma 3 27B ran on two NVIDIA L40S 48GB GPUs via vLLM [53] for faster inference. For Qwen 2 Audio, we used two NVIDIA GeForce RTX 3090 for LID and topic classification, and a Quadro RTX 6000 & Tesla V100S-PCIE-32GB for S2TT, RC-QA, and ASR.
@@ V Results And Discussions
Africa 0.0 56.6 68.5 28.2 55.3 68.5 26.3 45.6 55.5 83.9 60.1 18.0 13.2 28.5 37.6 18.2 52.5 64.2
Americas/Oceania 0.3 98.1 96.7 15.2 66.9 44.7 27.1 45.6 50.9 146.2 31.7 19.5 11.9 36.5 27.2 14.2 70.2 64.1
Asia (S) 4.5 91.8 97.5 32.8 79.8 83.0 30.3 62.2 71.5 113.7 35.1 7.7 16.0 47.3 58.3 22.5 78.0 84.0
Asia (SE) 10.5 88.8 91.0 33.0 76.0 82.3 31.4 63.8 75.8 97.8 35.0 7.3 16.0 48.0 57.8 24.9 76.2 83.0
Asia (W, C) 21.7 80.2 88.0 28.4 79.1 82.1 30.1 64.6 74.7 111.7 27.2 6.7 15.6 47.5 55.5 26.7 74.6 81.6
Asia (E) 59.2 90.5 98.1 58.7 80.5 83.5 43.5 70.0 77.3 48.2 23.2 13.9 34.0 47.2 52.8 53.8 80.3 86.3
Europe (W, N, S) 28.6 86.1 84.5 54.5 84.0 82.1 41.6 74.3 81.7 62.6 12.4 7.3 29.6 56.9 58.3 41.5 81.5 82.8
Europe (E) 11.8 90.4 94.1 53.6 85.6 84.9 36.6 74.0 82.1 98.2 12.0 9.9 23.9 59.3 62.2 34.0 83.3 87.0
Average (macro) 17.1 85.3 89.8 38.0 75.9 76.4 33.4 62.5 71.2 95.3 29.6 11.3 20.0 46.4 51.2 29.5 74.6 79.1
Average (micro) 15.8 81.2 86.1 40.4 76.0 79.5 33.3 63.4 72.6 88.4 30.1 10.3 20.5 47.3 53.5
@@ V-A Overall Results
Figure 1: Average results by tasks. For the ASR, we reported “100CER100-CER” value.
@@ V-A1 Results by Task
Figure 1 shows the result of various LLMs and the best result obtained from a task-specific model (i.e. Task Baseline). For the speech tasks, we find that for classification tasks (i.e. LID, TC, and RC-QA), our result shows that the task baseline achieved much better result, for example, the MMS-LID accuracy is more than +10+10 points better than the best LLM (Gemini 2.0 Flash). Similarly, for TC and RC-QA, the performance gap is more than +8+8 and +13+13 points respectively. Overall, we saw a wide performance gap between Qwen 2 Audio and Gemini 2.0 Flash which clearly shows that the former only covers a few languages. For more challenging tasks such as ASR and S2TT, surprisingly, we found Gemini 2.0 Flash to be better than SeamlessM4T v2, one of the SOTA task-specific baseline for both tasks. This result shows that Gemini 2.0 Flash should be preferred for these speech tasks especially for low-resource languages rather than their task-specific variants. For the text-based tasks, the finding is different, we find that the gap in performance between an open weight model such as Gemma 3 27B and Gemini 2.0 Flash or GPT-4o is much smaller. Similarly, the performance of task baselines are also very similar to those of the LLMs on almost all tasks except for LID and RC-QA. For LID, task baselines (GlotLID) is better by +5+5 points accuracy when compared to Gemini 2.0 Flash, while for RC-QA, all LLMs are significantly better than task baseline based on Translate-TRAIN (on XLM-V) with over +17+17 points.
@@ V-A2 Results by Region
Table  III and IV shows the result across eight different geographical groups. For the speech tasks, we find that the African and the Americas/Oceania region achieved the worst result across all the covered regions probably because most of the low-resource languages are concentrated there. While Qwen 2 Audio model achieved the worst results among the LLMs we evaluated, we find that its performance is slightly better for Asia (E) and Europe (W, N, S) which may reflect where most of the models pre-training data are coming from. For example on ASR and S2TT, it achieved less than 50 CER and more than 30 ChrF on East Asia region. Gemini 2.0 Flash on the other hand had less than 15 CER for all regions except Americas/Oceania and Africa. However, GPT-4o was worse than Gemini 2.0 Flash across all regions we tested, it only achieved less than 15 CER for the European regions. The worst performance of GPT-4o is also for African languages at 60.1 CER. We observe similar trend for S2TT where Gemini 2.0 Flash achieved more than 50 ChrF++ on all regions except the Americas/Oceania and African regions. For classification tasks, Gemini 2.0 Flash had over 70 accuracy points on TC and RC-QA except for the African and Americas region; however, for the LID, the most language confusion is for the African region. For text-based tasks, results were generally high except in Africa. While we observe the Americas region also has high performance across tasks including MT, this is not true for NLI where it achieved the lowest ($46.8$), more than 17-17 points behind Africa. A possible reason is that large datasets like Flores and SIB-200 cover few Americas/Oceania languages (about five, including two Creole languages), whereas AmericasNLI includes more truly low-resource ones (e.g., Otomi, Wixarika), offering a more realistic picture of performance of indigenous Americas languages.
@@ V-A3 Results by Language Family
Text-based tasks achieve overall higher performance: In general, performance on texts is higher than performance on speech for same tasks. This implies that current LLMs are better on texts than speech, which implies that more investment in developing better audio LLMs are needed. On the LID task, several language families especially those that use a different script have perfect accuracy score using Gemini 2.0 Flash, however, we noticed some mis-classifications for the spoken utterances. Similar trend is observed for TC and RC-QA classification tasks; on TC, all language families except Nilotic achieved more than $80$ accuracy point. Also, for the translation tasks, several language families achieved impressive ChrF++ of $60$ points and above in xx-en direction while only one language family (Uralic) achieved up to 60 ChrF++ in S2TT.
Language families in Africa achieved the lowest performance: For both text and speech modalities, the language families with the lowest performance are: Atlantic-Congo, Nilotic, and Afro-Asiatic, many or all of the languages in the first two language families are often categorized as low-resourced according to Joshi’s classification [57] and are only located in Africa. This confirms our findings in Sec. V-A2. Apart from having low performance, the gap in performance is also quite wide. For example, the average ChrF++ for S2TT in Nilotic and Atlantic-Congo is 36.5-36.5 and 25.9-25.9 lower when compared to Indo-European languages. Similarly, for RC-QA, the performance of Indo-European is $78.4$ while Atlantic-Congo, Nilotic, and Afro-Asiatic achieved 36.5-36.5, 25.9-25.9, and 25.9-25.9 respectively.
@@ V-B Detailed analysis of task specific results
The Global NLI dataset has the highest representation of Americas/Oceania amongst all the other datasets that we evaluated, with 11 out of the 59 languages being from this region. This is due to the Americas NLI dataset [30] and only 3 languages of this region are present in the other text-based tasks. On this task, GPT-4o performs the best overall and in most regions. Americas/Oceania is the region with the lowest performance in this task. We see a performance gap of 35-35 points between Americas/Oceania (49.1%) and Europe (W, N, S) (84.8%), which is larger than that present in other tasks. The NLI results show that many Americas/Oceanian languages are less covered by many LLMs, although only few multilingual datasets includes these indigeneous languages. More efforts are needed to expand benchmarks for this region of languages.
@@ Acknowledgment
This research was supported by the Google grant via Mila and the IVADO R3AI Regroupement Grant. Luel acknowledges the support of Carnegie Corporation of New York provided through AIMS-RIC. We would also like to thank Google Cloud for the GCP credits Award through the Gemma 2 Academic Program, and OpenAI for providing us access for providing API credits.